# Coarse-Grained Games: A Framework for Bounded Perception in Game Theory

### v1

@@ Abstract
In everyday life, we frequently make coarse-grained judgments. When we say that Olivia and Noah excel in mathematics, we disregard the specific differences in their mathematical abilities. Similarly, when we claim that a particular automobile manufacturer produces high-quality cars, we overlook the minor variations among individual vehicles. These coarse-grained assessments are distinct from erroneous or deceptive judgments, such as those resulting from student cheating or false advertising by corporations. Despite the prevalence of such judgments, little attention has been given to their underlying mathematical structure. In this paper, we introduce the concept of coarse-graining into game theory, analyzing games in which players may perceive different payoffs as identical while preserving the underlying order structure. This framework allows us to examine the rational inference processes that arise when players equate distinct micro-level payoffs at a macro level, and to explore how Nash equilibria are preserved or altered as a result. Our key findings suggest that coarse-grained games possess several desirable properties that make them suitable for modeling phenomena in the social sciences. This paper demonstrates two such applications: first, in cases of overly minor product updates, consumers may encounter an equilibrium selection problem, resulting in market behavior that is not driven by objective quality differences; second, the lemon market can be analyzed not only through objective information asymmetry but also through asymmetries in perceptual resolution or recognition ability.
@@ 1.1 Background
Game theory has seen significant developments since its formalization by [18]. The Folk Theorem, one of the fundamental results in repeated game theory, states that in infinitely repeated games, any feasible and individually rational payoff can be sustained as a Nash equilibrium [4], provided that players are sufficiently patient. This theorem assumes that players can accurately perceive differences in payoffs and respond accordingly.
@@
By applying a similar conceptual framework, we can analyze how coarse-graining in human decision-making allows individuals and institutions to manage complexity by filtering out fine details that are not operationally significant at a broader scale. To facilitate such an analysis, we seek to elucidate the mathematical structure of coarse-grained games, providing a formal framework for understanding how strategic interactions evolve when players perceive payoffs at different levels of granularity.
@@ 1.2 Comparison to Previous Research
The following section explains how coarse-grained games differ from existing game-theoretic frameworks:
@@
  • Coarse-grained games are not a subclass of imperfect information games. In a coarse-grained game, players observe their opponent’s strategies and payoffs through their own level of coarse perception. Since players are not assumed to be aware of differences in each other’s levels of coarse perception, they subjectively perceive the game as one of perfect information. This differs from imperfect information games, where it is explicitly assumed that certain strategies or payoffs remain unobservable to the players [19].
@@
  • Coarse-grained games are not Bayesian games. In a Bayesian game [8, 9, 10], each player possesses a probability distribution over types assigned by nature. In contrast, players in a coarse-grained game do not operate with such probabilistic beliefs but instead perceive the game through a coarse-grained payoff matrix. In this structure, payoffs are not represented as precise real numbers but as sets of real values, making them fundamentally different from probability distributions. Note: Depending on the choice of strategy preprocessing methods as discussed in this paper, a coarse-grained game could potentially be reformulated as a Bayesian game. However, this paper does not explore that possibility.
@@
  • Coarse-Grained Game is not a deception game [24]. The fact that one player perceives information with a different level of granularity is not because they have been misled by someone else, nor is it the result of an opponent deliberately providing false information. For example, when a teacher tells a student, “Your report was graded as excellent," instead of saying, “Your report received a score of 93,” they are neither lying nor spreading misinformation. The distinction lies in the coarseness of the information provided, not in its accuracy or truthfulness.
@@
  • Coarse-Grained Game does not concern naive players. In [3] framework, naive players are those who misaccount for selection, leading to incorrect decision-making. However, in a Coarse-Grained Game, the fundamental distinction lies between high-resolution players and low-resolution players, as will be discussed later. While low-resolution players perceive payoffs with less clarity, this does not imply that they are naive. On the contrary, a low-resolution player can sometimes be the smarter choice in a practical sense. For example, instead of a teacher who meticulously distinguishes between a student who scored 91 and another who scored 92, adjusting their grading method to reflect this one-point difference, a teacher who simply categorizes both as “excellent” may be seen as more efficient and judicious.
@@
  • A coarse-grained game is related to information asymmetry but does not necessarily presuppose it. As shown in Section 7.1, even if the player who initially possesses more information voluntarily and actively shares it to eliminate the asymmetry, similar phenomena may still arise if the other party lacks the sensitivity to properly interpret that information. In this case, the asymmetry lies not in the amount of information available, but in the precision with which it is perceived. Therefore, coarse-grained games cannot simply be reduced to a framework of information asymmetry; rather, they exhibit the characteristics of an epistemic framework.
@@ 1.3 Structure of This Paper
The structure of this paper is as follows: Section 2 defines coarse-grained games from the perspective of coarse set theory, establishing a formal foundation for their analysis. Section 3 interprets the meaning of payoffs in coarse-grained games, introducing the distinction between subjective and objective payoffs. Section 4 examines how Nash equilibria transform in coarse-grained games, focusing on one-shot (non-repeated) strategic interactions. Section 5 explores how changes in Nash equilibria can lead to unexpected incidental gains or losses for players in coarse-grained games, highlighting the potential for unforeseen strategic outcomes. Section 6 analyzes infinitely repeated games, demonstrating that variations in perceived discount factors may prevent cooperation at discount rates that players would otherwise consider sufficient in coarse-grained payoff matrices. Section 7 explores the applicability of coarse-grained games in the social sciences by analyzing how consumers’ coarse-grained perceptions influence decision-making. To illustrate this, we examine two cases: minor model changes in product markets and the lemon market. Section 8 concludes the paper with a summary of key findings and directions for future research.
@@ A Coarse-Grained Game $\mathcal{G}$ is defined as a tuple:
  • $\Phi=(\phi_{1},\phi_{2},\dots,\phi_{k},\dots,\phi_{n})$ is the tuple of coarse-graining functions, where each $\mathfrak{G}_{k}$ is assigned to player $k$:
    $\phi_{k}:\mathbb{R}\to\mathfrak{G}_{k};$
    $\quad\quad\phi_{k}(x)=G,\quad\text{where }x\in G\text{ and }G\in\mathfrak{G}_{k}.$
@@
The function $\Phi$ is a tuple of coarse-graining functions, where each $\phi_{k}$ applies a player-specific transformation to payoffs. If the function is to be applied to the entire matrix rather than to individual payoffs, it is acceptable to write $\Phi_{k}(M)$, provided that no confusion arises. $\Phi_{k}$ maps the detailed payoffs to a coarser representation based on the assigned partition $\mathfrak{G}_{k}$:
@@
In the following, we refer to the payoff matrix constructed to apply $\phi_{k}$ as $M^{\text{base}}$ which is determined by the combination of $N$, $S$, and $u$. The coarse-grained payoff matrix obtained by player $k$ through this application is denoted as $M^{\mathfrak{G}_{k}}_{k}$. Since the subscription of a coarse-grained payoff matrix is equal to the subscription of $\mathfrak{G}$, we omit the former as $M^{\mathfrak{G}_{k}}$. Applying $\psi_{k}$ to $M^{\mathfrak{G}_{k}}$ yields the final payoff matrix of each player, which we denote as $M^{\prime}_{k}$.
@@ 2.5.2 Example: Coarse-Grained Prisoners Dilemma
Applying this partition to the base payoff matrix $M^{\text{base}}$ through $\phi_{1}$:
@@
Applying this partition to $M^{\text{base}}$ through $\phi_{2}$:
@@ 3 Interpreting Coarse-Grained Payoffs
  • Objective Payoff within $M^{\text{base}}$: The base payoff matrix represents the true underlying structure of the game. This matrix is independent of player perception and reflects the game’s fundamental incentives. However, in a coarse-grained game, players do not have direct access to $M^{\text{base}}$.
@@ 4.1 Existence of Nash Equilibrium
Before exploring these considerations in depth, we must first analyze the characteristics of Nash equilibria in the coarse-grained game. Specifically, we will demonstrate that, even in such games, a Nash equilibrium is guaranteed to exist under appropriate conditions.
@@
Let $\mathcal{G}=(N,S,u,\mathfrak{G},\Phi,\Psi)$ be a coarse-grained game with mixed strategies and the EMP. Then, $\mathcal{G}$ has at least one mixed strategy Nash equilibrium in each $M^{\prime}_{k}$.
@@ 4.2 Preservation of the Original Nash Equilibrium with Pure Strategies
Next, we are interested in how the Nash equilibrium of the base payoff matrix changes for each player in the coarse-grained game. The following theorem establishes a result regarding this question.
@@
In a coarse-grained game that adopts EMP, the base payoff matrix undergoes a two-step transformation. First, it is mapped by $\Phi$ into a coarse-grained payoff matrix, where each cell contains a totally ordered set of real values rather than a single real number, aligning with each player’s resolution level. Then, each cell is further transformed by $\Psi$, which applies the aforementioned equation to convert it back into a real-valued payoff.
@@
$\quad\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots s^{*}_{n})))<\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}))).$
@@
Since the base set of $\mathfrak{G}$ is $(\mathbb{R},\leq)$, it is guaranteed that such $G\in\mathfrak{G}_{k}$ and $G^{\prime}\in\mathfrak{G}_{k}$ always exist in the coarse-grained game, and they are given by $\phi_{k}$ mapping $\mathbb{R}$ into $\mathfrak{G}$ as follows:
$G=\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n}))\text{ and }G^{\prime}=\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}),$
@@
$\iff\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n}))\succcurlyeq\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n})).$
@@
$\iff\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n})))\geq\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}))).$
@@
Theorem 3 states only that a pure strategy Nash equilibrium, rationally computed in the base matrix, remains a pure strategy Nash equilibrium in the coarse-grained game under certain conditions. It does not imply that a strategy profile that was not a pure strategy Nash equilibrium in the base matrix will necessarily remain non-equilibrium in the coarse-grained game. This becomes evident when considering a coarse-grained partition $\mathfrak{G}$ that includes a set $G$ containing all real values appearing in the base payoff matrix.
@@ 4.3.1 Case Study
In a coarse-grained game, it has already been shown that if mixed strategies and EMP are allowed, at least one Nash equilibrium exists. However, is this Nash equilibrium identical to that of the non-coarse-grained $M^{\text{base}}$? Unlike in the case of pure strategies, the mixed strategy Nash equilibrium in the original payoff matrix is not necessarily preserved as a mixed strategy Nash equilibrium in the coarse-grained game. This can be seen through the following simple counterexample.
@@
Applying $\phi_{1}$ and $\psi_{\text{EMP}}$ to the original matrix, we obtain $M^{\prime}_{1}$:
@@ 4.3.2 Uniform Strategy Reduction
$\quad\quad\forall s_{k},s_{-k},s^{\prime}_{k},s^{\prime}_{-k},\quad\psi_{\text{EMP}}(\phi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k}))),$
@@
Applying $\phi_{1}$ and $\psi_{\text{EMP}}$ to the original matrix, we obtain $M^{\prime}_{1}$:
@@
The function $\phi_{l}$ is a mapping that replaces a real-valued number $x$ with the corresponding set $G$ that contains $x$. In other words,
$\phi_{l}(x)=G,\quad\text{where }x\in G.$
@@
$u_{k}(s_{k},s_{-k})\in G\implies\phi_{l}(u_{k}(s_{k},s_{-k}))=\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})).$
@@
$\phi_{l}(u_{k}(s_{k},s_{-k}))=\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})),$
@@
$\psi_{\text{EMP}}(\phi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k}))).$
@@
$\forall s_{k},s_{-k},s^{\prime}_{k},s^{\prime}_{-k},\quad\psi_{\text{EMP}}(\phi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})))$
@@
$\phi_{l}(u_{k}(s_{k},s_{-k}))=\phi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})).$

### v2

@@ Abstract
In everyday life, we frequently make coarse-grained judgments. When we say that Olivia and Noah excel in mathematics, we disregard the specific differences in their mathematical abilities. Similarly, when we claim that a particular automobile manufacturer produces high-quality cars, we overlook the minor variations among individual vehicles. These coarse-grained assessments are distinct from erroneous or deceptive judgments, such as those resulting from student cheating or false advertising by corporations. Despite the prevalence of such judgments, little attention has been given to their underlying mathematical structure. In this paper, we introduce the concept of coarse-graining into game theory, analyzing games where players may perceive different payoffs as identical while preserving the underlying order structure. We call it a Coarse-Grained Game (CGG). This framework allows us to examine the rational inference processes that arise when players equate distinct micro-level payoffs at a macro level, and to explore how Nash equilibria are preserved or altered as a result. Our key findings suggest that CGGs possess several desirable properties that make them suitable for modeling phenomena in the social sciences. This paper demonstrates two such applications: first, in cases of overly minor product updates, consumers may encounter an equilibrium selection problem, resulting in market behavior that is not driven by objective quality differences; second, the lemon market can be analyzed not only through objective information asymmetry but also through asymmetries in perceptual resolution or recognition ability.
@@ 1.1 Background
Game theory has seen significant developments since its formalization by [19]. The Folk Theorem, one of the fundamental results in repeated game theory, states that in infinitely repeated games, any feasible and individually rational payoff can be sustained as a Nash equilibrium [5, 1], provided that players are sufficiently patient and that effective punishment strategies are available to deter deviations. This theorem assumes that players can accurately perceive differences in payoffs and respond accordingly.
@@
By applying a similar conceptual framework, we can analyze how coarse-graining in human decision-making allows individuals and institutions to manage complexity by filtering out fine details that are not operationally significant at a broader scale. To facilitate such an analysis, we seek to elucidate the mathematical structure of coarse-grained games (CGGs), providing a formal framework for understanding how strategic interactions evolve when players perceive payoffs at different levels of granularity.
@@ 1.2 Comparison to Previous Research
The following list explains how CGGs differ from existing game-theoretic frameworks:
@@
  • CGGs are not a subclass of imperfect information games. In a CGG, players observe their opponent’s strategies and payoffs through their own level of coarse perception. Since players are not assumed to be aware of differences in each other’s levels of coarse perception, they subjectively perceive the game as one of perfect information. This differs from imperfect information games, where it is explicitly assumed that certain strategies or payoffs remain unobservable to the players [20].
@@
  • CGGs are not Bayesian games. In a Bayesian game [9, 10, 11], each player possesses a probability distribution over types assigned by nature. In contrast, players in a CGG do not operate with such probabilistic beliefs but instead perceive the game through a coarse-grained payoff matrix (Section 2.2). In this structure, payoffs are not represented as precise real numbers but as sets of real values, making them fundamentally different from probability distributions. Note: Depending on the choice of strategy preprocessing methods as discussed in this paper (Section 2.3), a CGG could potentially be reformulated as a Bayesian game. However, this paper does not explore that possibility.
@@
  • CGG is not a deception game [25]. The fact that one player perceives information with a different level of granularity is not because they have been misled by someone else, nor is it the result of an opponent deliberately providing false information. For example, when a teacher tells a student, “Your report was graded as excellent," instead of saying, “Your report received a score of 93,” they are neither lying nor spreading misinformation. The distinction lies in the coarseness of the information provided, not in its accuracy or truthfulness.
@@
  • CGG does not concern naive players. In [4] framework, naive players are those who fail to account for selection, leading to incorrect decision-making. However, in a CGG, the fundamental distinction lies between high-resolution players and low-resolution players, as will be discussed later (Section 2.4). While low-resolution players perceive payoffs with less clarity, this does not imply that they are naive. On the contrary, a low-resolution player can sometimes be the smarter choice in a practical sense. For example, instead of a teacher who meticulously distinguishes between a student who scored 91 and another who scored 92, adjusting their grading method to reflect this one-point difference, a teacher who simply categorizes both as “excellent” may be seen as more efficient and judicious.
@@
  • CGG is related to information asymmetry but does not necessarily presuppose it. As shown in Section 7.1, even if the player who initially possesses more information voluntarily and actively shares it to eliminate the asymmetry, similar phenomena may still arise if the other party lacks the sensitivity to properly interpret that information. In this case, the asymmetry lies not in the amount of information available, but in the precision with which it is perceived. Therefore, CGGs cannot simply be reduced to a framework of information asymmetry; rather, they exhibit the characteristics of an epistemic framework.
@@ 1.3 Structure of This Paper
The structure of this paper is as follows: Section 2 defines CGGs from the perspective of coarse set theory, establishing a formal foundation for their analysis. Section 3 interprets the meaning of payoffs in CGGs, introducing the distinction between subjective and objective payoffs. Section 4 examines how Nash equilibria transform in CGGs, focusing on one-shot (non-repeated) strategic interactions. Section 5 explores how changes in Nash equilibria can lead to unexpected incidental gains or losses for players in CGGs, highlighting the potential for unforeseen strategic outcomes. Section 6 analyzes infinitely repeated games, demonstrating that variations in perceived discount factors may prevent cooperation at discount rates that players would otherwise consider sufficient in coarse-grained payoff matrices. Section 7 explores the applicability of CGGs in the social sciences by analyzing how consumers’ coarse-grained perceptions influence decision-making. To illustrate this, we examine two cases: minor model changes in product markets and the lemon market. Section 8 concludes the paper with a summary of key findings and directions for future research.
@@ A Coarse-Grained Game $\mathcal{G}$ is defined as a tuple:
  • $\Phi=(\varphi_{1},\varphi_{2},\dots,\varphi_{k},\dots,\varphi_{n})$ is the tuple of coarse-graining functions, where each $\mathfrak{G}_{k}$ is assigned to player $k$:
    $\varphi_{k}:\mathbb{R}\to\mathfrak{G}_{k};$
    $\quad\quad\varphi_{k}(x)=G,\quad\text{where }x\in G\text{ and }G\in\mathfrak{G}_{k}.$
@@
The function $\Phi$ is a tuple of coarse-graining functions, where each $\varphi_{k}$ applies a player-specific transformation to payoffs. If the function is to be applied to the entire matrix rather than to individual payoffs, it is acceptable to write $\Phi_{k}(M)$, provided that no confusion arises. $\Phi_{k}$ maps the detailed payoffs to a coarser representation based on the assigned partition $\mathfrak{G}_{k}$:
@@
In the following, we refer to the payoff matrix constructed to apply $\varphi_{k}$ as $M^{\text{base}}$ which is determined by the combination of $N$, $S$, and $u$. The coarse-grained payoff matrix obtained by player $k$ through this application is denoted as $M^{\mathfrak{G}_{k}}_{k}$. Since the subscription of a coarse-grained payoff matrix is equal to the subscription of $\mathfrak{G}$, we omit the former as $M^{\mathfrak{G}_{k}}$. Applying $\psi_{k}$ to $M^{\mathfrak{G}_{k}}$ yields the final payoff matrix of each player, which we denote as $M^{\prime}_{k}$.
@@ 2.5.2 Example: Coarse-Grained Prisoners Dilemma
Applying this partition to the base payoff matrix $M^{\text{base}}$ through $\varphi_{1}$:
@@
Applying this partition to $M^{\text{base}}$ through $\varphi_{2}$:
@@ 3 Interpreting Coarse-Grained Payoffs
  • Objective Payoff within $M^{\text{base}}$: The base payoff matrix represents the true underlying structure of the game. This matrix is independent of player perception and reflects the game’s fundamental incentives. However, in a CGG, players do not have direct access to $M^{\text{base}}$.
@@ 4.1 Existence of Nash Equilibrium
Before exploring these considerations in depth, we must first analyze the characteristics of Nash equilibria in the CGG. Specifically, we will demonstrate that, even in such games, a Nash equilibrium is guaranteed to exist under appropriate conditions.
@@
Let $\mathcal{G}=(N,S,u,\mathfrak{G},\Phi,\Psi)$ be a CGG with mixed strategies and the EMP. Then, $\mathcal{G}$ has at least one mixed strategy Nash equilibrium in each $M^{\prime}_{k}$.
@@ 4.2 Preservation of the Original Nash Equilibrium with Pure Strategies
Next, we are interested in how the Nash equilibrium of the base payoff matrix changes for each player in the CGG. The following theorem establishes a result regarding this question.
@@
In a CGG that adopts EMP, the base payoff matrix undergoes a two-step transformation. First, it is mapped by $\Phi$ into a coarse-grained payoff matrix, where each cell contains a totally ordered set of real values rather than a single real number, aligning with each player’s resolution level. Then, each cell is further transformed by $\Psi$, which applies the aforementioned equation to convert it back into a real-valued payoff.
@@
$\quad\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots s^{*}_{n})))<\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}))).$
@@
Since the base set of $\mathfrak{G}$ is $(\mathbb{R},\leq)$, it is guaranteed that such $G\in\mathfrak{G}_{k}$ and $G^{\prime}\in\mathfrak{G}_{k}$ always exist in the CGG, and they are given by $\varphi_{k}$ mapping $\mathbb{R}$ into $\mathfrak{G}$ as follows:
$G=\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n}))\text{ and }G^{\prime}=\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}),$
@@
$\iff\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n}))\succcurlyeq\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n})).$
@@
$\iff\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s^{*}_{k},\dots,s^{*}_{n})))\geq\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{*}_{1},s^{*}_{2},\dots,s_{j_{k}},\dots,s^{*}_{n}))).$
@@
Theorem 3 states only that a pure strategy Nash equilibrium, rationally computed in the base matrix, remains a pure strategy Nash equilibrium in the CGG under certain conditions. It does not imply that a strategy profile that was not a pure strategy Nash equilibrium in the base matrix will necessarily remain non-equilibrium in the CGG. This becomes evident when considering a coarse-grained partition $\mathfrak{G}$ that includes a set $G$ containing all real values appearing in the base payoff matrix.
@@ 4.3.1 Case Study
In a CGG, it has already been shown that if mixed strategies and EMP are allowed, at least one Nash equilibrium exists. However, is this Nash equilibrium identical to that of the non-coarse-grained $M^{\text{base}}$? Unlike in the case of pure strategies, the mixed strategy Nash equilibrium in the original payoff matrix is not necessarily preserved as a mixed strategy Nash equilibrium in the CGG. This can be seen through the following simple counterexample.
@@
Applying $\varphi_{1}$ and $\psi_{\text{EMP}}$ to the original matrix, we obtain $M^{\prime}_{1}$:
@@ 4.3.2 Uniform Strategy Reduction
$\quad\quad\forall s_{k},s_{-k},s^{\prime}_{k},s^{\prime}_{-k},\quad\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k}))),$
@@
Applying $\varphi_{1}$ and $\psi_{\text{EMP}}$ to the original matrix, we obtain $M^{\prime}_{1}$:
@@
The function $\varphi_{l}$ is a mapping that replaces a real-valued number $x$ with the corresponding set $G$ that contains $x$. In other words,
$\varphi_{l}(x)=G,\quad\text{where }x\in G.$
@@
$u_{k}(s_{k},s_{-k})\in G\implies\varphi_{l}(u_{k}(s_{k},s_{-k}))=\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})).$
@@
$\varphi_{l}(u_{k}(s_{k},s_{-k}))=\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})),$
@@
$\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k}))).$
@@
$\forall s_{k},s_{-k},s^{\prime}_{k},s^{\prime}_{-k},\quad\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s_{k},s_{-k})))=\psi_{\text{EMP}}(\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})))$
@@
$\varphi_{l}(u_{k}(s_{k},s_{-k}))=\varphi_{l}(u_{k}(s^{\prime}_{k},s^{\prime}_{-k})).$